OpenStudy (anonymous):

Evaluate f(x) = 4x + 3 for x = -2

OpenStudy (anonymous):

-5

OpenStudy (rulnick):

f(-2) = 4(-2) + 3 = -8 + 3 = -5

OpenStudy (anonymous):

You like put -2 in for x, and then you, like, do algebra and stuff.